Operation Summary for September 4, 2009 - September 7, 2009 |
Friday September 4 Startup. Beam studies in progress for MiniBooNE. Operations reported having database problems; controls personnel were investigating. During the day shift, operators prepared the Main Injector (MI) for access; MiniBooNE beam was off during the morning for this access. Tevatron (TeV) personnel conducted a ground fault search and found the problem. Controls personnel continued their investigation of a vacuum problem. During the evening shift, the West Anode Power Supply tripped off due to a Booster RF station (BRF18). Operators bypassed the station and brought BRF19 on in its place. The Pelletron was conditioning in cooling mode and the TeV separators were also conditioning. Saturday September 5 Startup. MiniBooNE continued its beam studies. During the midnight shift, a Linac Klystron RF station (KRF3) tripped off; it reset. BRF17 tripped off; it reset. During the day shift, operators prepared MI for access to check on the bake. MiniBooNE resumed taking beam around 11:05 AM. Controls personnel investigated a cold-cathode vacuum readback problem at TeV sector B3. During the evening shift, the TeV cooldown was started at 5:50 PM. Sunday September 6 Startup. MiniBooNE continued to take beam when it was available. During the midnight shift, controls personnel fixed the TeV sector B3 vacuum readback problem. A Booster power supply (B:ORBUMP) tripped off; operators found and replaced a bad fuse. During the day shift, operators prepared MI for a short access. KRF3 tripped off several times; it reset. The TeV cooldown continued and the Antiproton Source (Pbar) and MI bakes continued. During the evening shift, the Linac low energy Buncher tripped off; it reset. Monday September 7 Startup. MiniBooNE continued to take beam when it was available. During the midnight shift, the MI and Pbar bake-outs continued as well the TeV cooldown. The Plans for Monday and for the rest of the Week Startup work continues. Pbar expects their bake to finish later today. The MI bake seems to be going slower than expected. MiniBooNE would like to begin optic studies on Tuesday, 9/8/09. Cryo system personnel believe that the TeV cooldown will end on Tuesday night, 9/8/09. |
|
More Information Comments and Suggestions |
| last modified 09/08/2009 email Fermilab |
| Security, Privacy, Legal | |